OpenStudy (anonymous):

For each of the following sets of Vectors, determine whether it is a basis in R3. If a set is a basis find the coordiantes of a = (0,1,3) in this basis. a) (2)(0) (-3)(5) (4)(6) b) (1)(0)(1) (0)(1)(1) (-1)(2)(1) c) (1)(0)(1) (0)(1)(0) (-1)(2)(1)

OpenStudy (anonymous):

Do you know what a basis is?

OpenStudy (anonymous):

Isn't a basis a set of one non zero vector v in the direction of l if it is for a line?

OpenStudy (anonymous):

That's true for a line or R1, how about R3?

OpenStudy (anonymous):

is it a set of 3 non-coplanar vectors?

OpenStudy (anonymous):

That's right. So are the vectors from a) a basis?

OpenStudy (anonymous):

Is it a set of three vectors?

OpenStudy (anonymous):

Yeah it is a set of 3 vectors isn't it?

OpenStudy (anonymous):

O wait no

OpenStudy (anonymous):

A is only 2 sets where is B and C contain 3 sets right?

OpenStudy (anonymous):

A is a set of two vectors and B an C are sets of three vectors.

OpenStudy (anonymous):

Yeah that's what I mean :) What do I do after to figure out if it is a basis of a = 1,3,0...can I draw it out or is there a way to do it through an equation

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

If the set is a basis, it asks to find the coordinates of a which is (0,1,3) how do I do that?

OpenStudy (anonymous):

Right, have you learned how to solve systems of equations? With Matrices?

OpenStudy (anonymous):

No we haven't yet

OpenStudy (anonymous):

Too bad, it's a lot easier if you had. Anyway let's do it for c). Finding coordinates means that you're looking for constants a, b and c, such that a*(1,0,-1)+b(0,1,2)+c(1,0,1)=(0,1,3).

OpenStudy (anonymous):

You can see right away that b is equal to 1, because (1,0,-1) and (1,0,1) both have zeros on the second place.

OpenStudy (anonymous):

What remains is: a+b*0+c=0 and a*-1+2b+c=3

OpenStudy (anonymous):

So a+c=0 and c-a=1. So a=-c, thus c+c=1, so c=1/2 and a =-1/2

OpenStudy (anonymous):

how do you get a+c=0 and c-a = 1? I got lost there

OpenStudy (anonymous):

o nvm

OpenStudy (anonymous):

you just simplified the equations since B = 0

OpenStudy (anonymous):

I made a mistake on that second one, sorry about that.

OpenStudy (anonymous):

Where did you make a mistake?

OpenStudy (anonymous):

oh you mean c-a=3?

OpenStudy (anonymous):

No it was actually fine before.

OpenStudy (anonymous):

Sorry about this confusion.

OpenStudy (anonymous):

a*-1+2b+c=3 -a+2+c=3 c-a=1

OpenStudy (anonymous):

How is 2 within the equation I thought B=0?

OpenStudy (anonymous):

b is equal to one.

OpenStudy (anonymous):

Oh yeah! This is making sense now. What you do with a+c=o and c-a=1 after?

OpenStudy (anonymous):

How did you get c+c=1?

OpenStudy (anonymous):

from a+c=0 it follows that c=-a Now I substitute c for -a in c-a=1, so you get c+c=1

OpenStudy (anonymous):

AHHHH THIS MAKES SO MUCH SENSE YOU ARE THE BEST :) Could I ask you one more thing?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Could you explain the idea of Linear Independence and how to figure out if a set spans linearly independent?

OpenStudy (anonymous):

Do you know the dot product or inner product of two vectors?

OpenStudy (anonymous):

Well she is just teaching us that but I don't think we have to apply dot product to our assignment because she isn't even finished teaching us it.

OpenStudy (anonymous):

Two vectors are independent if there not on the same line. So (1,0,0) and (2,0,0) are dependent, while (1,2,0) and (1,0,5) are independent.

OpenStudy (anonymous):

A vector can also be independent from a set of vectors. This is the case when that vector lies outside the plane spanned by the set. For instance (1,0,0) is independent from the set {(0,1,0),(0,01)}. And (0,2,3) is dependent from the set {(0,1,0),(0,0,1)}

OpenStudy (anonymous):

So a(-1)(2) and b(1)(3) are linearly independent right?

OpenStudy (anonymous):

What do you mean by a and b?

OpenStudy (anonymous):

Well it asks if these sets are linearly independent which are a) (1) (-1) (2) (-1) (2) (3) b) (-1) (1) (2) (3)

OpenStudy (anonymous):

A set of vectors can also be linearly independent, this is the case all the vectors are linearly independent from the rest of the vectors in the set. (So a and b are seperate questions.) A set is dependent if you can take a combination of the vectors equal to one. For instance if you have the set {v,w,u}, and a*v+b*w+c*u=0 for some numbers a,b c (not all zero) then it is dependent.

OpenStudy (anonymous):

Alright thank you :) I am stuck on the Question where we have to figure out the coordinates of A. I have been trying to do B and I got to this a+c=0 b+c=1 -a+2b+c=3 I have tried many numbers but it doesn't work? Can you guide me once again <3

OpenStudy (anonymous):

Well, it's no wonder you can't find them, because B is not a basis. It's not an independent set.

OpenStudy (anonymous):

so it isn't possible to do?

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

Sweet Thank you! You have been awesome! Thanks for guiding me through my confusing of Linear Algebra :)

OpenStudy (anonymous):

You're welcome. Keep at it, you'll develop a feeling for it and it'll become a lot easier.